OpenStudy (anonymous):

How do I factor 64 + 48t -16t^2=0?

OpenStudy (anonymous):

-4t(4t-12-16)=0?

OpenStudy (anonymous):

correct..

OpenStudy (anonymous):

to factor this further.....

OpenStudy (anonymous):

-4t(4t-12-16)=0 -16t^2+48t+64t wrongg!!

OpenStudy (anonymous):

64 + 48t -16t^2=0 16(4+3t-t^2)=0

OpenStudy (anonymous):

16(4+3t-t^2)=0 is answer!!!

OpenStudy (anonymous):

ok @malibugranprix2000

OpenStudy (mayankdevnani):

right..